OpenStudy (anonymous):

covert 21/30 to a percebt plz or a t least a bit of help on how too xD im not the biggest math person there is

OpenStudy (anonymous):

Note that \(\dfrac{21}{30} =\dfrac{7\cdot 3}{10\cdot 3} = \dfrac{7}{10}\). Do you think you can write this as a percentage now? :-)